# Generalized Pinching-Antenna Systems: A Tutorial on Principles, Design Strategies, and Future Directions

### v1

@@ V-A Pinching-Antenna System Design under LoS Blockage
Looking forward, several open challenges merit further investigation. These include the development of hybrid channel models that account for both LoS and NLoS propagation in the context of pinching-antenna topologies, robust optimization algorithms that are resilient to partial or uncertain blockage, and real-time reconfiguration schemes that adapt to dynamic environments. Addressing these issues will be key to unlocking the full potential of pinching-antenna systems in next-generation wireless networks operating under severe propagation constraints.
@@ V-D Accurate and Realistic Electromagnetic Modeling of Pinching Antennas
V-E Pinching-Antenna-Enabled Extremely Large-Scale Antenna Array
@@
Extremely large-scale antenna array (ELAA) systems are emerging as a key technology to meet the ambitious performance demands of next-generation wireless networks [4, 134]. By dramatically increasing the number of antennas, ranging from hundreds to thousands, ELAA systems deliver significant improvements in spatial resolution, spectral efficiency, link reliability, and energy efficiency. These capabilities translate into better wireless channel sounding, more precise beamforming, and the ability to serve massive numbers of users simultaneously. ELAA can be realized through two primary approaches: (i) single-BS deployment, where a BS aggregates a vast array of antennas into a centralized structure; and (ii) coordinated distributed deployment, where multiple spatially separated access points, each equipped with a subset of antennas, are jointly coordinated to function as a virtual super-array. Both forms of ELAA offer substantial gains but also present critical implementation challenges, including high interconnection costs, computational complexity, and practical implementation challenges, e.g., synchronization and backhaul limitations for distributed ELAA [4, 135, 136].
Pinching antennas present a highly promising enabler for practical ELAA deployment. Their unique capability to flexibly reconfigure antenna positions along dielectric waveguides enables not only compact integration of massive arrays in a single BS, but also the rapid and low-cost expansion of distributed antenna networks. By leveraging the strong LoS links and reduced path loss inherent to pinching-antenna architectures, ELAA systems can be deployed with enhanced channel quality and spatial selectivity, while also reducing hardware and cabling complexity. For single-BS ELAA, pinching antennas facilitate the dense packing and flexible placement of a large numbers of elements, supporting both near-field and far-field beamforming. In distributed ELAA scenarios, pinching antennas allow for dynamic assignment and reconfiguration of antennas across access points, enabling adaptive user-centric coverage, simplified maintenance, and more robust scalability.
@@
Going forward, the integration of pinching antennas into ELAA architectures opens several exciting research directions. These include the development of distributed signal processing and beamforming algorithms tailored for flexible arrays, robust calibration and synchronization techniques in dynamically reconfigurable environments, and the joint optimization of pinching-antenna positioning with resource allocation and user association strategies. Overall, pinching-antenna-enabled ELAA systems are well positioned to address many of the fundamental challenges in future wireless networks, and represent a promising pathway toward practical, scalable, and high-performance ELAA deployments.
V-F Pinching Antenna-Assisted Wireless Edge Intelligence
@@
Edge AI is envisioned as a key enabler for next-generation wireless networks [137, 138, 139, 140]. By bringing data processing and model training closer to end user devices, wireless edge intelligence supports a wide range of intelligent applications such as federated learning, autonomous driving, smart manufacturing, and real-time environment perception. However, the performance of distributed AI algorithms in wireless networks is often limited by imperfect and dynamically changing propagation conditions, such as fading and interference [141, 142, 143, 144]. These impairments can significantly degrade the convergence, reliability, and latency of wireless model training and inference.
@@
V-G Robust Pinching-Antenna System Design
The performance of pinching-antenna systems is intrinsically linked to accurate user position information, even under LoS channel conditions. In practical deployments, however, the precise user locations are often unavailable due to limitations in positioning technologies, user mobility, and measurement errors [146, 147]. These uncertainties can lead to considerable performance degradation, as antenna placement and beamforming strategies become suboptimal in the presence of position errors. Despite its practical significance, the robust design of pinching-antenna systems under user position uncertainty remains largely unexplored in the current literature.
Addressing this open challenge will require the development of robust optimization frameworks that explicitly incorporate user position uncertainty into system design. Potential approaches may include maximizing worst-case performance metrics (such as the minimum data rate or SNR), minimizing outage probability, or leveraging stochastic models for uncertainty [148, 149]. These problems are inherently complex, often demanding new algorithmic strategies that jointly optimize antenna positioning and resource allocation while maintaining computational efficiency for dynamic and large-scale networks. Future research in this area is both necessary and promising. Effective robust pinching-antenna system designs could significantly improve the reliability and consistency of next-generation wireless networks, especially in dynamic or high-mobility environments.
V-H Real-Time Low-Complexity Control of Mechanical Antenna Placement
@@ VI Conclusion
Pinching-antenna systems offer a flexible and innovative solution to many of the limitations faced by conventional antenna technologies in next-generation wireless networks. By enabling dynamic antenna placement and activation along dielectric waveguides, pinching-antenna systems provide enhanced LoS connectivity, spatial adaptability, and strengthened support for advanced functionalities such as integrated sensing and security. In this tutorial, we have developed a generalized pinching-antenna framework, where different mediums, such as dielectric waveguides and LCX, can be used with different antenna activation methods. Then, we have introduced the underlying physical mechanisms and distinctive channel characteristics of representative pinching-antenna system architectures, encompassing configurations with both single and multiple waveguides, and discussed state-of-the-art system design strategies. We also reviewed the integration of pinching antennas with emerging wireless technologies and highlighted a broad range of promising application scenarios. Despite these significant advances, practical deployment of pinching-antenna systems continues to face several challenges. Notable issues include the need for robust designs that account for user position uncertainty, the development of real-time, low-complexity hardware and control solutions, and deeper integration with distributed intelligence and edge computing paradigms. Addressing these open problems will be essential for fully realizing the benefits of pinching-antenna systems in future wireless environments. In summary, while substantial progress has been achieved, continued research and innovation are imperative to unlock the full potential of pinching-antenna systems and to ensure their successful adoption in the evolving landscape of wireless communications.

### v2

@@ V-A Pinching-Antenna System Design under Random LoS/NLoS Channels
Initial studies on pinching-antenna system design under random LoS/NLoS channels in [94] have already shown that pinching antennas can provide significant gains in such challenging propagation environments, see Section III-C. However, several open challenges merit further investigation, including scalable designs for multi-waveguide pinching-antenna topologies, robust optimization algorithms that are resilient to partial or uncertain blockage, and real-time reconfiguration schemes that can adapt to highly dynamic environments. Addressing these issues will be key to fully unlocking the potential of pinching-antenna systems in next-generation wireless networks operating under severe propagation constraints.
@@ V-D Accurate and Realistic Electromagnetic Modeling of Pinching Antennas
V-E Environment-Aware Pinching-Antenna System Designs
Environment-aware designs, which consider factors such as user data traffic patterns and obstacle patterns, are crucial for optimizing coverage, capacity, and reliability in dynamic and complex environments. This is particularly attractive for outdoor or highly dynamic scenarios where per-user tracking is difficult. Notably, environment-aware designs typically facilitate long-term network planning, as the environment generally does not change rapidly. Such designs allow networks to adapt to environmental conditions, ensuring robust performance even under challenging scenarios. However, traditional antenna systems are often limited by their fixed deployment, lacking the adaptability to respond to changing user distributions and evolving environmental factors.
Pinching-antenna systems are particularly well-suited for environment-aware designs due to their inherent flexibility. By adjusting the positions of antennas along pre-deployed waveguides, pinching antennas can dynamically respond to shifting user demands and changing environments. This adaptability makes them ideal for scenarios where conventional static antennas would perform poorly. Whether responding to hotspots in high-traffic areas or navigating obstacles in complex indoor environments, pinching-antenna systems offer a unique ability to optimize network performance through reconfigurable antenna placement over the long term.
Despite their potential, several challenges remain for realizing the full benefits of environment-aware pinching-antenna systems. Key challenges include developing efficient optimization algorithms for dynamic antenna placement, integrating both traffic-aware and geometry-aware models, and ensuring system robustness under fluctuating environmental conditions. Overcoming these challenges will be crucial for unlocking the full potential of pinching-antenna systems in real-world, long-term network deployments.
V-F Pinching-Antenna-Enabled Extremely Large-Scale Antenna Array
@@
Extremely large-scale antenna array (ELAA) systems are emerging as a key technology to meet the ambitious performance demands of next-generation wireless networks [7, 144]. By dramatically increasing the number of antennas, ranging from hundreds to thousands, ELAA systems can substantially improve spatial resolution, spectral efficiency, link reliability, and energy efficiency. ELAA can be realized through two primary approaches: (i) single-BS deployment, where a BS aggregates a vast array of antennas into a centralized structure; and (ii) coordinated distributed deployment, where multiple spatially separated access points, each equipped with a subset of antennas, are jointly coordinated to function as a virtual super-array. While both forms offer significant gains, they also face critical implementation challenges, including high interconnection costs, computational complexity, and practical issues such as synchronization and backhaul limitations [7, 145, 146].
Pinching antennas present a promising enabler for practical ELAA deployment. In a typical architecture, multiple waveguides are deployed and each waveguide hosts a large number of pinching antennas along its length. This allows a very large number of radiating points to be realized from a moderate number of feeder points and RF chains, enabling compact massive arrays at a single BS and low-cost expansion of distributed antenna networks. A recent study on large-scale multi-cell deployment of pinching antennas [147] further confirms the potential of such architectures, showing that appropriately designed waveguide-based deployments can reshape cell boundaries, improve worst-case cell performance, and enhance network-wide data rates. By leveraging the strong LoS links and reduced path loss inherent to pinching-antenna architectures, such ELAA systems can offer enhanced channel quality and spatial selectivity while significantly reducing hardware and cabling complexity. In both single-BS and distributed ELAA scenarios, multiple waveguides with reconfigurable pinching antennas support scalable multi-user services via spatial multiplexing across waveguides and conventional resource-domain schemes (e.g., OMA/NOMA).
@@
Going forward, the integration of pinching antennas into ELAA architectures opens several exciting research directions. These include the design of distributed signal processing and beamforming algorithms tailored to flexible pinching-antenna arrays, the development of robust calibration and synchronization techniques in dynamically reconfigurable environments, and the joint optimization of pinching-antenna positioning with resource allocation and user association. Overall, pinching-antenna-enabled ELAA systems offer a promising pathway toward practical, scalable, and cost-effective ELAA deployments.
V-G Pinching Antenna-Assisted Wireless Edge Intelligence
@@
Edge AI is envisioned as a key enabler for next-generation wireless networks [148, 149, 150, 151, 152]. By bringing data processing and model training closer to end user devices, wireless edge intelligence supports a wide range of intelligent applications such as federated learning, autonomous driving, smart manufacturing, and real-time environment perception. However, the performance of distributed AI algorithms in wireless networks is often limited by imperfect and dynamically changing propagation conditions, such as fading and interference [153, 154, 155, 156]. These impairments can significantly degrade the convergence, reliability, and latency of wireless model training and inference.
@@
V-H Robust Pinching-Antenna System Design
The performance of pinching-antenna systems is intrinsically linked to accurate user position information and CSI. In practical deployments, however, the precise user locations and accurate CSI are often unavailable due to limitations in positioning technologies, user mobility, and measurement errors [158, 159]. These uncertainties, encompassing both user position errors and CSI uncertainty, can lead to significant performance degradation, as antenna placement, beamforming strategies, and resource allocation become suboptimal when errors are present.
Despite its practical importance, the robust design of pinching-antenna systems under user position and CSI uncertainty remains largely unexplored in the current literature. Addressing this open challenge will require the development of robust optimization frameworks that explicitly incorporate both user position uncertainty and CSI errors into the system design. Potential approaches may include maximizing worst-case performance metrics (such as the minimum data rate or SNR), minimizing outage probability, or leveraging stochastic models to represent these uncertainties [160, 161]. These problems are inherently complex, often demanding new algorithmic strategies that jointly optimize antenna positioning, resource allocation, and beamforming, while maintaining computational efficiency for dynamic and large-scale networks.
V-I Real-Time Low-Complexity Control of Mechanical Antenna Placement
@@
V-J Practical Implementation Challenges and Scalability
Beyond the algorithmic and architectural issues discussed above, the real-world deployment of generalized pinching-antenna systems raises several practical challenges that remain largely unexplored. First, fabrication tolerances of the guiding media and pinching elements (e.g., variations in dielectric profile, LCX slot geometry, connector quality, and mechanical alignment) can affect in-medium attenuation, coupling efficiency, and phase consistency along the waveguide or conductor, and thus need to be captured by more realistic models. Second, large-scale deployments with multiple waveguides and many pinching points will require efficient calibration of both amplitude and phase, potentially in near-field operation and dynamic repositioning of radiating spots. Finally, scalability becomes a key concern when the number of reconfigurable points grows large, since control signaling, channel estimation, and coordination overhead can easily become bottlenecks. Addressing these issues will require joint hardware–algorithm co-design, measurement-driven modeling, and the development of scalable calibration and control schemes tailored to generalized pinching-antenna architectures.
Complementing wired-fed architectures, recent work on wireless-fed pinching-antenna systems has opened up new possibilities for large-scale and multi-cell deployments [165]. By feeding the waveguides through relay antennas instead of dedicated wired links, wireless-fed pinching-antenna systems preserve the core advantages of pinching antennas (guided propagation plus flexible radiating spots) while alleviating the cost, fragility, and routing constraints of long end-to-end waveguide connections. This makes them particularly attractive for multi-cell scenarios, where multiple BSs or relay nodes can cooperatively illuminate large areas using networks of pinching-enabled waveguides. At the same time, such architectures introduce additional practical challenges, including the joint design of BS–relay–waveguide topologies, management of inter-cell interference, synchronization across wirelessly fed segments, and scalable control of many distributed pinching points. Systematically addressing these issues is an important step toward truly massive, network-wide deployments of generalized pinching-antenna systems.
@@ VI Conclusion
Pinching-antenna systems offer a flexible and innovative solution to many of the limitations faced by conventional antenna technologies in next-generation wireless networks. By enabling dynamic antenna placement and activation along dielectric waveguides, pinching-antenna systems provide enhanced LoS connectivity, spatial adaptability, and strengthened support for advanced functionalities such as integrated sensing and communications. In this paper, we have developed a generalized pinching-antenna framework, where different media, such as dielectric waveguides and LCX, different antenna activation methods can be employed. Then, we have introduced the underlying physical mechanisms and distinctive channel characteristics of representative pinching-antenna system architectures, encompassing configurations with both single and multiple waveguides, and discussed state-of-the-art system design strategies. We also reviewed the integration of pinching antennas with emerging wireless technologies and highlighted a broad range of promising application scenarios. Despite these significant advances, practical deployment of pinching-antenna systems continues to face several challenges. Notable issues include the need for robust designs that account for user position uncertainty, the development of real-time, low-complexity hardware and control solutions, and deeper integration with distributed intelligence and edge computing paradigms. Addressing these open problems will be essential for fully realizing the benefits of generalized pinching-antenna systems in future wireless environments. In summary, while substantial progress has been achieved, continued research and innovation are imperative to unlock the full potential of generalized pinching-antenna systems and to ensure their successful adoption in the evolving landscape of wireless communications.